OpenStudy (anonymous):

simplify (2√5+3√7)^2

OpenStudy (anonymous):

\[(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

no I meant like what are the steps

OpenStudy (jlastino):

@AldousJames dude he's not allowed to use calculators to find square roots of numbers

OpenStudy (anonymous):

\[(2\sqrt{5}+3\sqrt{7})(2\sqrt{5}+3\sqrt{7})\] \[2\sqrt{5}\times 2\sqrt{5}+2\times 2\sqrt{5}\times 3\sqrt{7}\times 3\sqrt{7}\times 3\sqrt{7}\]

OpenStudy (anonymous):

\[=4\times5+12\sqrt{35}+9\times 7\] \[=20+12\sqrt{35}+63\] \[=83+12\sqrt{35}\]

OpenStudy (anonymous):

83+ 12^0.5

OpenStudy (anonymous):

four multiplications just like \((a+b)^2=a^2+2ab+b^2\)

OpenStudy (anonymous):

The task was to simplify so and not to calculate so I guess using a calculator would be quite wrong. satellite73 did it right.